I cannot find my error.
I am supposed to use the quadratic formula to solve:
x^2+4x=5
This is what i got:
x=-4+-radical4^2-4*1*5/2*1
-4+-radical 16-20/2
-4+-radical -4/2
-4+-2i/2
-2+-i
Hope that makes sense. I am frustrated with this problem

Found 3 solutions by solver91311, checkley77, stanbon:
Answer by solver91311(24713)



Your error is that you did not put the quadratic into standard form first.

Standard form is:



But your



is actually in



form.

Either way, your 5 in the quadratic formula has to be -5, thus:



So:



or




Check:












John


Answer by checkley77(12844)

x^2+4x=5
x^2+4x-5=0 You forgot to transfer the 5.
You also need to use {( }) for this equation so you won't get lost.
This is what i got:
x=(-4+-radical[4^2-4*1*-5])/2*1
x=(-4+-radical[16+20])/2
x=(-4+-radical36)/2
x=(-4+-6)/2
x=(-4+6)/2
x=2/2=1 ans.
x=(-4-6)/2
x=-10/2
x=-5 ans.

Answer by stanbon(75887)

use the quadratic formula to solve:
x^2+4x=5
Rearrange:
x^2 + 4x -5 = 0
----
x=[-4+-radical(4^2-4*1*-5)]/(2*1)
x = [-4+-radical(16+20)]/2
x = [-4+-radical(36)]/2
x = [-4 +- 6]/2
x = [-10/2] or x = [2/2]
x = -5 or x = 1
=====================
Cheers,
Stan H.